OpenStudy (anonymous):

\[\int_0^{\infty } 2^t e^{-s t} \, dt\]

OpenStudy (anonymous):

2^t= e^Log[2^t]= e^2Log[2] \[\int_0^{\infty } e^{-s t} e^{t \log (2)} \, dt\] \[\int _0^{\infty } e^{t \log (2)-s t}dt\]

OpenStudy (cwrw238):

is this = 1 / (log2 - s) * e ^ (tlog2 - st) im not very confident in integration

OpenStudy (anonymous):

just \[\frac{1}{s-\text{Log}[2]}\]

OpenStudy (cwrw238):

- this is the result after applying the limits?

OpenStudy (cwrw238):

your first line should be = e^tLog[2] right?

OpenStudy (cwrw238):

turning 2^t to an exponential function is pretty clever!

OpenStudy (anonymous):

yes, after limit

OpenStudy (cwrw238):

gret

OpenStudy (anonymous):

cwrw238 are you jimmyrep? :-)